OpenStudy (anonymous):

solve -1/6m=22 132 22/6 -132 -22/6

OpenStudy (callisto):

Divide both sides by -1/6, what do you get?

OpenStudy (anonymous):

idk.

OpenStudy (callisto):

Alright, multiply both sides by -6, that do you get?

OpenStudy (callisto):

*what do you get

OpenStudy (anonymous):

3.666

OpenStudy (anonymous):

-

OpenStudy (callisto):

\[-\frac{1}{6}m=22\]Multiply both sides by -6 \[-6(-\frac{1}{6}m)=-6(22)\]\[m=-6(22)=...?\]